OpenStudy (anonymous):

(x+1)(x+2)(x+3)(x+4) = 99 (by hand)

OpenStudy (anonymous):

\[t=x^2+5x\]

OpenStudy (shubhamsrg):

that was pure genius !

OpenStudy (anonymous):

ty :)

OpenStudy (anonymous):

Enlighten me....

OpenStudy (anonymous):

Or better, tell me what t is?

OpenStudy (shubhamsrg):

(x+1)(x+4)(x+2)(x+3) = 99 (x^2 + 5x +4)(x^2 + 5x + 6) = 99 => (x^2 + 5x)^2 + 10 (x^2 + 5x) + 75 = 0 easier after that..

OpenStudy (anonymous):

(x^2 + 5x +4)(x^2 + 5x + 6) = 99 (y-1)(y+1) = 99

OpenStudy (shubhamsrg):

aha,,cool !

OpenStudy (anonymous):

ya much simpler :)

OpenStudy (anonymous):

Yours is also a timesaver:-)

OpenStudy (anonymous):

It will work whenever there is a sequence (arithmetic) or some symmetry (like 1,5,7,11)